Check whether or not * is a binary operation on S={x∈R| x>0}, where x*y=| In(xy)|


1
Expert's answer
2021-03-01T07:10:20-0500

By definition of a binary operation, we should have for any "x,y\in S", "x*y\in S". But if we take "x\in S, y=\frac{1}{x}" (that always exists, as "x\neq 0"), we obtain "x*y=|\ln(x\cdot \frac{1}{x})|=|0|\notin S". Thus the image of "*" is not contained in "S", so "*" is not a binary operation.